# Simulator-Free Visual Domain Randomization via Video Games

### v1

@@ 2 Background
Video Games for CV. Inspired by insights from Risi & Togelius (2020) and Reed et al. (2022) suggesting that procedurally generated sets of diverse games enhance generality in machine learning, we explore the use of existing video games in CV. Several recent studies investigate the use of commercial-standard games as an alternative to dedicated in-lab simulators or procedural game level generation approaches, in an attempt to circumvent limitations related to inaccessible game engines. Notably, Grand Theft Auto 5 (Rockstar, 2013) serves as a popular video game for collecting annotated data, achieved by intercepting rendering communication between graphics hardware and the screen buffer (Richter et al., 2016) or employing a game modification such as “infrared vision mod” (Gu et al., 2023). Alternatively, Pearce & Zhu (2022) gather internal game state information from CS:GO by probing the machine’s memory. In contrast to such prior works involving the “reverse-engineering” of game engines, our approach instead captures high-level game information such as player actions using raw inputs from the machine’s I/O devices, thereby simplifying the collection of annotated gameplay and boosting the scalability of our method across numerous commercial video games.
@@ 4.2 Behavior Alignment
Sensitivity Analysis of k: An insightful study from the perspective of domain randomization involves analyzing how the number of diverse games (k) in BehAVE’s training process affects alignment performance. For that purpose, we explore the impact of varying the number of games included in the training set, ranging from 1 to 15 and compare them on a fixed test set of 10 games. We randomize the selection of games in the train and test sets across the 5 independent runs reported in Section 5.1.
@@ 6 Discussion
Ethical Considerations. Recognizing inherent risks in FPS game data and ethical considerations in deploying such data for real-world applications is crucial. A comprehensive analysis of these implications is imperative for responsible use and future research in this area (Melhart et al., 2023).
@@ Appendix C Behavior Alignment
Algorithm 2 provides a PyTorch-like pseudocode implementation of the BehAVE framework depicted in Algorithm 1. Table 6 presents extended results from a comprehensive list of all configurations tested in the experiments reported in Section 5.1. Notably, in addition to varying the foundational video and text encoders, we also vary the alignment model architecture (see Section C.1) as well as the loss function employed in the behavior alignment training (see Section C.2.
@@ C.2 Varying the Loss Function
In contrast to the single-data-point focus of Cosine and MSE losses, we explore a pairwise comparison approach using the preference learning paradigm. This approach is motivated by the challenges discussed in Section A.2, where not every keypress corresponds to a visible animation, and vice versa. For example, a mouse left click action may not lead to a gun fire animation within the game if the weapon is empty, hence keypresses are only considered as “proxy-labels” for annotating on-screen animations. To accommodate this nuance in our recorded dataset, we adopt a preference learning method that avoids strict assumptions about keypress labels. Thus, the conditional preference loss in Equation 3 (with a margin hyper-parameter $\lambda$) can be conceptualized as a term that specifically considers the relative proximity of a video embedding to its corresponding text action embedding, when compared to the embedding of a different video. This formulation introduces a more lenient constraint on the alignment process, reducing susceptibility to training noise stemming from the “proxy-labels”.
$\mathcal{L}_{\text{pref}}(z^{\text{align}}_{i},z^{\text{align}}_{j}\mid z^{\text{caption}}_{i})=\max\left(0,\mathcal{L}_{\text{cos}}(z^{\text{align}}_{i},z^{\text{caption}}_{i})-\mathcal{L}_{\text{cos}}(z^{\text{align}}_{j},z^{\text{caption}}_{i})+\lambda\right)$ (3)
@@ D.1 Varying the Video Encoder
Table 8: Transferability of behavior classification from CS;GO (source) to SMG-25 test set (target). Actions having less than $30\%$ data points present in the training dataset are highlighted in gray.

### v2

@@ 2 Background
Video Games for CV. Inspired by insights from Risi & Togelius (2020) and Reed et al. (2022) suggesting that procedurally generated sets of diverse games enhance generality in machine learning, we explore the use of existing video games in CV. Several recent studies investigate the use of commercial-standard games as an alternative to dedicated in-lab simulators or procedural game level generation approaches, in an attempt to circumvent limitations related to inaccessible game engines. Notably, Grand Theft Auto 5 (Rockstar, 2013) serves as a popular video game for collecting annotated data, achieved by intercepting rendering communication between graphics hardware and the screen buffer (Richter et al., 2016; Krähenbühl, 2018; Taesiri et al., 2022) or employing a game modification such as “infrared vision mod” (Gu et al., 2023). Alternatively, Pearce & Zhu (2022) gather internal game state information from CS:GO by probing the machine’s memory. In contrast to such prior works involving the “reverse-engineering” of game engines, our approach instead captures high-level game information such as player actions using raw inputs from the machine’s I/O devices, thereby simplifying the collection of annotated gameplay and boosting the scalability of our method across numerous commercial video games.
@@ 4.2 Behavior Alignment
Sensitivity Analysis of k. An insightful study from the perspective of domain randomization involves analyzing how the number of diverse games (k) in BehAVE’s training process affects alignment performance. For that purpose, we explore the impact of varying the number of games included in the training set, ranging from 1 to 15 and compare them on a fixed test set of 10 games. We randomize the selection of games in the train and test sets across the 5 independent runs reported in Section 5.1.
@@ 7 Conclusion
8 Ethical Impact Statement
This paper introduces a dataset of fully anonymized annotated FPS gameplay videos. Both the gameplay videos and the annotations were collected in-house from participants in a laboratory setting under a data collection protocol approved by the University Research Ethics Committee of the University of Malta. The protocol was inspired by previously peer-reviewed work cited in Section II and is detailed in the Appendix A. We acknowledge the emerging trends in machine learning for autonomous weapons research (Simmons-Edler et al., 2024) and wish to clarify that this project has no military applications and was not funded by military sources. Additionally, the data collection protocol is exclusively applicable to games, and the dataset aims to support research in video games and artificial intelligence.
9 Dataset and Usage Rights
The dataset utilized in this research was derived from various commercial games. All game data consisting of screen captured game graphics, is the intellectual property of their respective game developers and publishers (please refer to Appendix A for proper attribution). The inclusion of the data in this study falls under the provisions of fair use for the purpose of academic research, analysis, and non-commercial study under US law, and text and data mining (TDM) for research purposes under EU law. The dataset will be made publicly available for future studies to ensure scientific reproducibility under a non-commercial license.
@@ Appendix C Behavior Alignment
Algorithm 2 provides a PyTorch-like pseudocode implementation of the BehAVE framework depicted in Algorithm 1. Table 6 presents extended results from a comprehensive list of all configurations tested in the experiments reported in Section 5.1. Notably, in addition to varying the foundational video and text encoders, we also vary the alignment model architecture (see Section C.1) as well as the loss function employed in the behavior alignment training (see Section C.2).
@@ C.2 Varying the Loss Function
In contrast to the single-data-point focus of Cosine and MSE losses, we explore a pairwise comparison approach using the preference learning paradigm. This approach is motivated by the challenges discussed in Section A.2, where not every keypress corresponds to a visible animation, and vice versa. For example, a mouse left click action may not lead to a gun fire animation within the game if the weapon is empty, hence keypresses are only considered as “proxy-labels” for annotating on-screen animations. To accommodate this nuance in our recorded dataset, we adopt a preference learning method that avoids strict assumptions about keypress labels. Thus, the conditional preference loss in Equation C.2 (with a margin hyper-parameter $\lambda$) can be conceptualized as a term that specifically considers the relative proximity of a video embedding to its corresponding text action embedding, when compared to the embedding of a different video. This formulation introduces a more lenient constraint on the alignment process, reducing susceptibility to training noise stemming from the “proxy-labels”.
{dmath}
L_pref(z^align_i, z^align_j ∣z^caption_i) = max(0, L_cos(z^align_i, z^caption_i) - L_cos(z^align_j, z^caption_i) + λ)
@@ D.1 Varying the Video Encoder
Table 8: Transferability of behavior classification from CS:GO (source) to SMG-25 test set (target). Actions having less than $30\%$ data points present in the training dataset are highlighted in gray.